how to solve x + 2y = - 1 and 2x - 3y = 12
Answer by sdmmadam@yahoo.com(530)

how to solve x + 2y = - 1 and 2x - 3y = 12
x + 2y = - 1 ----(1)This gives x =(-1-2y)----(*)(adding -2y to both the sides)
Putting (*) in 2x - 3y = 12 ----(2) we get
2(-1-2y)-3y = 12
-2-4y-3y = 12
-4y-3y = 2+12 (adding 2 to both the sides)
-7y =14
y = -14/7 = -2
Putting y = -2 in (*),
x=(-1-2y)= -1-2X(-2) = -1+4 = 3
Answer: x=3 and y = -2
Verification: x=3 and y = -2 in 2x - 3y = 12 ----(2)
LHS= 2x - 3y = 2X3-3X(-2) = 6+6= 12=RHS
Therefore our values are correct.